40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Dec 1 00:24:37 2012, Den, Ayaka, Update, Alignment, BS chamber push.png
    Reply  Sat Dec 1 13:56:38 2012, rana, Update, Alignment, BS chamber 
       Reply  Mon Dec 3 07:40:00 2012, Steve, HowTo, Alignment, how was the BS chamber misalinged BSsupport.jpg
Message ID: 7772     Entry time: Sat Dec 1 00:24:37 2012     Reply to this: 7773
Author: Den, Ayaka 
Type: Update 
Category: Alignment 
Subject: BS chamber 

Today at 11:13 AM the stack of invacuum BS table was kicked and IFO misaligned. We adjusted PZT2 voltage by ~20 V in yaw such that IPPOS was restored. Then we could lock arms.

push.png

ELOG V3.1.3-